# Enhancing Multistep Brent Oil Price Forecasting with a Multi-Aspect Metaheuristic Optimization Approach and Ensemble Deep Learning Models

### v1

@@ 5.1 Dataset
Figure 4: Brent price, USD index (USDX), and the sentiment score (SENT) for the period from 2012 to 2021
@@ 5.2 Experimental setup
As mentioned earlier, our experiments are conducted in two stages. In the first stage, the objective is to identify optimal values for hyperparameters of each model (optimal configuration). The targeted models are described in Section (3.3), and the targeted hyperparameters are described in Table (1). Subsequently, we train and evaluate each model individually using its corresponding optimal configuration. The forecasts obtained from each model serve as inputs for the second stage. In the second stage, we aim to optimally blend the forecasts obtained individually from each model in the first stage to produce the final forecast.
@@
time-steps (3-30) The number of timesteps used to represent the input data is crucial for capturing the temporal dynamics of the series.
@@
Generally, implementation of the GWO metaheuristic optimisation involves defining the main following procedures: problem preparation, objective function definition, population initialisation, and optimisation process. Firstly, the problem preparation involves specifying the hyperparameters to be optimised and establishing their respective ranges. The objective function is a pivotal element, representing the metric that the metaheuristic aims to either maximise or minimise during the optimisation process. This function typically encapsulates the performance of the deep learning model, such as accuracy or loss. The population in the context of metaheuristic optimisation refers to a collection of candidate solutions simultaneously. The metaheuristic explores and refines this population throughout the optimisation iterations to iteratively improve the model’s performance. A larger population generally provides more diversity and allows for a more thorough search space exploration, but it also increases computational cost. The hyperparameters and their respective ranges (bounds or search space) are as shown in Table (1).
@@
5.3 Objective function and evaluation metrics
For each model, the GWO has been run multiple independent runs with an objective function to minimise the Mean square error (MSE), which can be calculated according to the equation (5). However, to evaluate the final ensemble and individual models, we used the traditional ML measurements: mean square error (MSE), root mean square error (RMSE) - equation (6), and mean absolute error (MAE)- equation (7):
$\text{MSE}(y,\hat{y})=\frac{\sum_{i=0}^{N-1}(y_{i}-\hat{y}_{i})^{2}}{N}$
$\text{RMSE}(y,\hat{y})=\sqrt{\frac{\sum_{i=0}^{N-1}(y_{i}-\hat{y}_{i})^{2}}{N}}$
$\text{MAE}(y,\hat{y})=(\frac{1}{n})\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$
where $y$ indicates to the actual value, while $\hat{y}$ is the predicted value.
5.4 Weighted ensemble model
Inspired by the advantages of blending forecasts using conditional methods, we examine a meta-heuristic strategy for blinding the forecasts obtained from individual models. The weighted ensemble method in forecasting involves a mathematical approach to combine the predictions of multiple individual models to improve overall accuracy and robustness. Each model in the ensemble contributes to the final forecast with a certain weight. The final ensemble forecast is then obtained by summing the products of each model’s forecast with its corresponding optimal weight. Mathematically, this can be represented as the weighted sum:
Weighted ensemble prediction $=\sum_{i=1}^{n}M_{i}W_{i}$ (8)
Here, $W_{i}$ represents an optimal weight assigned to $model_{i}$, where $0\leq W_{i}\leq 1$, and $\sum W_{i}=1$. While $M_{i}$ represents $model_{i}$ forecasts.
In this experiment, a dynamic weighted ensemble method was employed to blend the predictions of the top five forecasts obtained from the SENT-Bi-GRU, SENT-Bi-LSTM, SENT-CNN-Bi-LSTM, SENT-CNN-Bi-LSTM-att, and SENT-USD-encoder-decoder-LSTM. The weights metric $W$ was optimised using the Grey Wolf Optimization (GWO) metaheuristic algorithm. The experiments showed that this metaheuristic-calibration-based weighted ensemble method achieves better results than traditional method or other naive ensemble methods and that the ensemble model outperforms the individual models, achieving a mean square error of (0.000165), with a success ratio of reducing the MSE 0.024%, compared to the best-performing individual model as shown in Table 4.
@@ 6 Results and discussion
In this section, we first present the optimal configuration obtained for each model identified by the GWO algorithm. Then, we discuss the accuracy of the forecast achieved by blending the best forecasts and evaluate the performance of the ensemble model by comparing the MAE, MSE and RMSE values to the individual model’s selected benchmark models.
@@ 6.1 Models optimal configurations
Table 2 shows the best solution found by GWO for each selected model. The first column (MSE) shows the best MSE achieved by each model, while the other columns show the corresponding optimal configuration values. The results obtained from the GWO-optimised five benchmark models, in Table 2, showed that the best-performing model is based on the Bi-GRU architecture when incorporating the sentimental score (SENT-Bi-GRU), with the best MSE of 0.000137. The corresponding optimal configuration identified in this model is AdamW as the optimiser with a sliding window of 5 lagged steps, 0.003154571 as the learning rate, 0.0399269013 as the dropout value, and 2 hidden units. Integrating the sentimental score index (SENT) generally gives better results than integrating the USDX index. While combining both factors improved the results only with the encoder-decoder-LSTM architecture. The optimal values for sliding windows are (5, 6, 10, 17). These figures align consistently with typical workdays, as 5 corresponds to the number of working days a week.
@@
6.2 GWO-ensemble model accuracy evaluation
@@
As shown in the table, the GWO-ensemble model outperformed the other models, achieving a 0.000127 of MSE. The best final forecast has been acquired by blending the forecasts obtained from SENT-Bi-GRU, SENT-Bi-LSTM, SENT-Bi-CNN-LSTM-attention, SENT-encoder-decoder-LSTM and SENT-USDX-encoder-decoder-LSTM models. The optimal weights assigned for each model are illustrated in Table (3).
@@
Model MAE MSE RMSE
GWO-Ensemble 0.007414 0.000127 0.011269
SENT-BI-GRU 0.008300 0.000137 0.011704
SENT-BI-LSTM 0.009800 0.000194 0.013928
SENT-CNN-BI-LSTM 0.012650 0.000280 0.016733
SENT-CNN-BI-LSTM-att 0.013550 0.000300 0.017306
SENT-USD-CNN-BI-LSTM-att 0.014400 0.000349 0.018682
SENT-Encoder-Decoder-BI-LSTM 0.012550 0.000359 0.018934
SENT-USD-Encoder-Decoder-BI-LSTM 0.012900 0.000344 0.018547
Figure 5 illustrates the forecasting performances for all targeted models. The figure shows the plot chart of Actual vs. predicted results. Table (4) shows the overall comparison of the GWO-ensemble model with the other benchmark models.
Figure 6: Objective function and run time charts of a single GWO run for the following models: (a,b):LSTM, (c,d): GRU, (e,f): CNN-LSTM-att, (g,h): encoder-decoder
@@ 8 Conclusions
In this study, we have introduced a hybrid approach that combines deep learning networks with metaheuristic optimisation to enhance the forecasting of Brent crude oil prices. The main idea is to harness the capabilities of metaheuristic optimisation to identify optimal configurations at various stages of constructing a forecasting solution, alongside the power of combining forecasts with multiple deep learning models. The well-known Grey Wolf Optimizer (GWO) algorithm was employed to optimise feature selection, dataset preparation, tuning hyperparameters for five individual deep learning networks, and the fusion of their forecasts. The results demonstrated that this approach improved the accuracy for each model and that the presented ensemble model outperforms benchmark models and achieves an MSE of 0.000127.

### v2

@@ 5.1 Dataset and input features
Figure 4: Historical data of Brent oil price, US dollar index (USDX), and the sentimental score (SENT)
@@ 5.2 Experimental setup
As mentioned earlier, our experiments are conducted in two stages. In the first stage, the objective is to identify optimal values for hyperparameters of each model (optimal configuration). The individual models are described in Section 4.1, and the targeted hyperparameters are described in Table 1. Subsequently, we train and evaluate each model individually using its corresponding optimal configuration. The forecasts obtained from each model serve as inputs for the second stage. In the second stage, we aim to optimally blend the forecasts obtained individually from each model in the first stage to produce the final forecast as illustrated in Figure 2.
@@
sliding window size (3-30) The number of timesteps used to represent the input data is crucial for capturing the temporal dynamics of the series.
@@
The hyperparameters and their respective ranges (bounds or search space) are as shown in Table 1.
@@
5.3 Evaluation metrics
For each model, the GWO has been run multiple independent runs with an objective function to minimise the Mean square error (MSE), which can be calculated according to the equation (6). However, to evaluate the final ensemble and individual models, we used the following six common evaluation metrics:
  • Mean Absolute Error (MAE) measures the average absolute difference between the predicted and actual values.
    $\text{MAE}(y,\hat{y})=\frac{\sum_{i}^{n}|y_{i}-\hat{y}_{i}|}{n}$ (5)
  • Mean Squared Error (MSE) for calculating the average squared difference between the predicted and actual values.
    $\text{MSE}(y,\hat{y})=\frac{\sum_{i}^{n}(y_{i}-\hat{y}_{i})^{2}}{N}$ (6)
  • Root Mean Squared Error (RMSE), providing a measure of the average magnitude of the error.
    $\text{RMSE}(y,\hat{y})=\sqrt{\frac{\sum_{i}^{n}(y_{i}-\hat{y}_{i})^{2}}{n}}$ (7)
  • Mean Squared Prediction Error (MSPE) is a statistical measure used to evaluate the accuracy of a predictive model. It quantifies the average squared difference between the predicted values and the actual observed values.
    $MSPE(y,\hat{y})=\frac{1}{n}\sum_{i}^{n}(y_{i}-\hat{y}_{i})^{2}$ (8)
  • Coefficient of Determination ($R^{2}$ or R-squared) is a statistical measure of how well the regression predictions approximate the actual data points. It is given by the following equation:
    $R^{2}(y,\hat{y})=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (9)
These evaluation metrics allow for quantifying the accuracy and precision of the model’s predictions. Lower MSE, MAE, MSPE, and RMSE values indicate better performance, While a high $R^{2}$ value indicates better performance.
@@ 6 Results and discussion
In this section, we first present the optimal configurations identified by the GWO algorithm for each model. We then compare the accuracy of the ensemble forecasts with that of the individual model forecasts.
@@ 6.1 Models optimal configurations
Table 2 shows the best solutions identified by GWO for each model. The first column shows the MSE value achieved by the model using the corresponding configuration values, as determined for each model by the GWO optimizer, which are displayed in the subsequent columns.
The results obtained from the GWO-optimised five benchmark models, in Table 2, showed that the best-performing model is based on the Bi-GRU architecture when incorporating the sentimental score (SENT-Bi-GRU), with the best MSE of 0.000137. The corresponding optimal configuration identified in this model is AdamW as the optimiser with a sliding window of 5 lagged steps, 0.003154571 as the learning rate, 0.0399269013 as the dropout value, and 2 hidden units. Integrating the sentiment score index (SENT) generally provides greater accuracy compared to the USDX index. However, combining both factors improves performance exclusively with the encoder-decoder-LSTM architecture. The optimal values for sliding windows are (5, 6, 10, 17). These figures align consistently with typical workdays, as 5 corresponds to the number of working days a week.
@@
6.2 GWO-ensemble model evaluation
@@
The experiments showed that this metaheuristic-calibration-based weighted ensemble method achieves better results than traditional method or other naive ensemble methods and that the ensemble model outperforms the individual models, achieving a mean square error of (0.000127) as shown in Table 4. The best final forecast has been acquired by blending the forecasts obtained from SENT-Bi-GRU, SENT-Bi-LSTM, SENT-Bi-CNN-LSTM-attention, SENT-encoder-decoder-LSTM and SENT-USDX-encoder-decoder-LSTM models. The optimal weights assigned for each model are illustrated in Table 3.
@@
Model MAE MSE RMSE MSPE MAPE $R^{2}$
GWO-Ensemble 0.007414 0.000127 0.011269 0.001649 0.026401 0.981353
SENT-BI-GRU 0.008300 0.000137 0.011704 0.002089 0.031694 0.980519
SENT-BI-LSTM 0.009800 0.000194 0.013928 0.002722 0.036012 0.972414
SENT-CNN-BI-LSTM 0.012650 0.000280 0.016733 0.001633 0.031219 0.853016
SENT-CNN-BI-LSTM-att 0.013550 0.000300 0.017306 0.004602 0.050919 0.957414
SENT-USD-CNN-BI-LSTM-att 0.014400 0.000349 0.018682 0.006660 0.057875 0.950375
SENT-Encoder-Decoder-BI-LSTM 0.012550 0.000359 0.018934 0.005175 0.046804 0.949024
SENT-USD-Encoder-Decoder-BI-LSTM 0.012900 0.000344 0.018547 0.005353 0.049209 0.951086
Figure 5 illustrates the forecasting performances for all individual models. The figure shows the plot chart of Actual vs. predicted results. Table 4 shows the overall comparison of the GWO-ensemble model with the other benchmark models.
Figure 6: Objective function and runtime charts of a single GWO run for various models: (a,b): LSTM, (c,d): GRU, (e,f): CNN-LSTM-att, (g,h): Encoder-Decoder.
@@ 8 Conclusions
In this study, we demonstrated how the accuracy of Brent oil price forecasting can be enhanced through the use of ensemble models and advanced optimization techniques. We have introduced a hybrid approach that combines established deep learning networks with Grey Wolf Optimizer (GWO) to enhance the forecasting of Brent crude oil prices. The main idea is to optimise configurations at various stages of constructing a forecasting solution, alongside the power of combining forecasts with multiple deep learning models. The GWO algorithm was employed to optimise feature selection, dataset preparation, tuning hyperparameters for five individual deep learning networks, and the fusion of their forecasts. The results demonstrated that this approach improved the accuracy for each model and that the presented ensemble model outperforms benchmark models and achieves an MSE of 0.000127.
Declarations
  • Funding: No external funding was received for this research.
  • Conflict of interest/Competing interests: Not applicable.
  • Ethics approval and consent to participate: Not applicable.
  • Data availability: The dataset used in this paper can be accessed through this GitHub repos
  • Code availability: Will be available upon request.
  • Corresponding author: Mohammed Alruqimi